1. The corporate bond matures in 20 years an is callable in 10, its current market value is $1035. ITs coupon rate is 5% paid annually. What is the yield-to-call. The call premium is $25.

2. What is the theoretical value of an investment that promises to pay you $75 at the end of every year for eight eras. Plus $4,000 at the end of the eight year. Your required rate of return is 6%.
Answer #1

1.Price =1035
Number of Years =20
Number of years =10
Coupon =5%*1000 =50
Future Value =Par Value+Call premium =1000+25 =1025

Yield To call using excel formula =RATE(10,50,-1035,1025)=4.75%

2. Price of Investment =PV of payments for 8 year + PV of 4000 at end of 8th year
=75*((1-(1+6%)^-8)/6%)+4000/(1+6%)^8 =2975.38